# Classification of locality preserving symmetries on spin chains

### v1

@@ E.2 Covariant right restrictions
for all $g,k\in G$. The failure of the left restriction to be covariant is captured by local unitaries $\operatorname{\Psi}_{g}(k)$ which are uniquely defined up to phase by
@@
for all $g,k,l,m\in G$. They therefore define a cohomology class $[\lambda]\in H^{2}(G,U(1)[G])$. (See Appendix G for the relevant definitions.)
The class $[\lambda]$ depends only on the symmetry $\alpha$, so we obtain a well defined map
@@
In particular, the obstruction lifts to a homomorphism of monoids $\Lambda:(\operatorname{\mathsf{Sym}}_{G}/\sim)\rightarrow H^{2}(G,U(1)[G])$.
Proof.
Can be found in Appendix I. ∎
@@
    We will show in Appendix J that the obstruction $\Lambda(\alpha)$ is given by the inverse of the slant product of the anomaly of $\alpha$ (see Appendix G for definitions). In the context of SPTs the obstruction $\Lambda$ is therefore intimately related to the types and properties of the anyons supported by the gauged SPT [18], which are believed to be described by the Dijkgraaf-Witten TQFT [19] corresponding to the anomaly.
As the name suggests, if a symmetry $\alpha$ has vanishing obstruction to covariant right restrictions, then (after adding an ancilla) $\alpha$ indeed admits covariant right restrictions:
@@
Let $\alpha$ be a symmetry of range $R$ such that $\Lambda(\alpha)=[1]$. For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with a gauge degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain $\widetilde{\mathcal{A}}$. Then the symmetry $\tilde{\alpha}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}$, where $\rho_{\operatorname{\mathrm{reg}}}$ is the regular representation of $G$ on the gauge degree of freedom, admits a covariant right restriction at $j$ with defect size $5R$.
@@
$V_{g}^{*}\,\tilde{\alpha}^{(k)}\big{(}V_{\bar{k}gk}\big{)}$ $=\left(\sum_{l_{1}}\,\operatorname{\Psi}_{g}(l_{1})^{*}\otimes|l_{1}\rangle\langle l_{1}|\right)\times\left(\sum_{l_{2}}\,\tilde{\alpha}^{(k)}\big{(}\operatorname{\Psi}_{\bar{k}gk}(k_{2})\big{)}\otimes|kl_{2}\rangle\langle kl_{2}|\right)$
putting $l=l_{1}=hl_{2}$ and using Eq. (E.2) this becomes
@@
Now consider the right restriction $\tilde{\beta}_{\geq}$ with components $\tilde{\beta}_{\geq}^{(g)}=\mathrm{Ad}[V_{g}]\circ\tilde{\alpha}_{\geq}^{(g)}$. Then
@@
Suppose $\alpha$ has range $R$. We can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an extended spin chain $\widetilde{\mathcal{A}}$. By Lemma E.1 there is a right restriction $\tilde{\alpha}_{\geq}$ of $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ at $j$ with defect size $5R$ and such that $g\mapsto\tilde{\alpha}_{\geq}^{(g)}$ is a group homomorphism. We can therefore regard $\tilde{\alpha}_{\geq}$ as a symmetry.
Since $\tilde{\alpha}_{\geq}$ and $\alpha$ agree everywhere to the right of the site $5R$ it follows from local computability of the obstruction to covariant right restrictions (item 2 of Proposition E.2) that $\Lambda(\alpha)=\Lambda(\tilde{\alpha}_{\geq})$. But $\tilde{\alpha}_{\geq}$ agrees with $\mathrm{id}$ everywhere to the left of the site $-5R$, so $\Lambda(\tilde{\alpha}_{\geq})=\Lambda(\mathrm{id})=[1]$ by local computability. We conclude that $\Lambda(\alpha)=[1]$. ∎
E.3 Covariant right restrictions that are group homomorphisms
@@
Let $\alpha$ be a symmetry of range $R$ with trivial anomaly on a spin chain ${\mathcal{A}}$. For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with two gauge degrees of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})\otimes\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain ${\mathcal{A}}^{\prime\prime}$. There is an extension $\alpha=\alpha^{\prime\prime}\otimes\rho$ of $\alpha$ to the enlarged spin chain and a covariant right restriction $\beta_{\geq}$ of $\alpha^{\prime\prime}$ at $j$ with defect size $7R$ such that $g\mapsto\beta_{\geq}^{(g)}$ is a group homomorphism.
@@
For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with a gauge degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain ${\mathcal{A}}^{\prime}$. Let $\alpha^{\prime}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}$. Then Lemmas E.4 and E.5 imply that there is a covariant right restriction $\alpha^{\prime}_{\geq}$ of $\alpha^{\prime}$ at $j$ of defect size $5R$.
Since $\alpha^{\prime}$ is a symmetry of range $R$ with trivial anomaly we can repeat the construction in the proof of Lemma E.1 to obtain from $\alpha^{\prime}_{\geq}$ a new right restriction that is a group homomorphism. We point out here that this construction does not spoil the covariance.
To this end, let $\operatorname{\Phi}(g,h)\in{\mathcal{A}}_{[j-5R,j+7R]}$ be fusion operators associated with $\alpha^{\prime}_{\geq}$, with phases chosen such that
$\operatorname{\Phi}(f,g)\operatorname{\Phi}(fg,h)=(\alpha^{\prime}_{\geq})^{(f)}\big{(}\operatorname{\Phi}(g,h)\big{)}\operatorname{\Phi}(f,gh).$ (E.3)
Further enlarging the on-site algebra ${\mathcal{A}}^{\prime}_{j}$ by tensoring with another gauge degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$ we obtain an enlarged spin chain ${\mathcal{A}}^{\prime\prime}$. As in the proof of Lemma E.1 we define unitaries $V(g)=\sum_{k}\operatorname{\Phi}(g,k)\otimes|k\rangle\langle gk|$ supported on $[j-5R,j+6R]$ so that $\beta_{\geq}^{(g)}=\mathrm{Ad}[V(g)^{*}]\circ(\alpha^{\prime}_{\geq}\otimes\mathrm{id})$ defines a left restriction of $\alpha^{\prime}\otimes\mathrm{id}$ such that $g\mapsto\beta_{\geq}^{(g)}$ is a group homomorphism.
Covariance of $\alpha^{\prime}_{\geq}$ applied to the definition of the fusion operators results in
$(\alpha^{\prime})^{(k)}\big{(}\operatorname{\Phi}(g,h)\big{)}=\mu_{k}(g,h)\times\operatorname{\Phi}(kg\bar{k},kh\bar{k})$
for some phases $\mu_{k}(g,h)$ that satisfy
$\mu_{k}(lg\bar{l},lh\bar{l})\mu_{l}(g,h)=\mu_{kl}(g,h)\quad\text{and}\quad\mu_{k}(f,g)\mu_{k}(fg,h)=\mu_{k}(g,h)\mu_{k}(f,gh).$
The first of these relations follows from computing $(\alpha^{\prime})^{(kl)}\big{(}\operatorname{\Phi}(g,h)\big{)}$ in two different ways. The second relation, which says that the $\mu_{k}$ are 2-cocycles, follows from (E.3).
Using these relations one easily verifies that
$\rho_{\mu}^{(k)}(|f\rangle\langle g|)=\frac{\mu_{k}(g,\bar{f})}{\mu_{k}(\bar{f},f)\mu_{k}(g,f)}|kf\bar{k}\rangle\langle kg\bar{k}|$
defines a unitary $G$-action on the gauge degree of freedom. Taking $\alpha^{\prime\prime}=\alpha^{\prime}\otimes\rho_{\mu}$ we find
$(\alpha^{\prime\prime})^{(k)}\big{(}V(g)\big{)}=\sum_{f\in G}\,\frac{\mu_{k}(g,f)\mu_{k}(gf,\bar{f})}{\mu_{k}(\bar{f},f)\mu_{k}(gf,f)}\operatorname{\Phi}(kg\bar{k},kf\bar{k})\otimes|kf\bar{k}\rangle\langle kg\bar{k}kf\bar{k}|=V(kg\bar{k})$
where we used the 2-cocycle relation for $\mu_{k}$ on elements $(gf,\bar{f},f)$.
This implies that $\beta_{\geq}^{(g)}=\mathrm{Ad}[V(g)^{*}]\circ(\alpha^{\prime}_{\geq}\otimes\mathrm{id})$ is a covariant right restriction of $\alpha^{\prime\prime}$ at $j$ with defect size $7R$. ∎
Appendix F Proof of the main Theorem
Proposition F.1.
@@ Appendix F Proof of the main Theorem
We will decouple the symmetry by splitting $\alpha$ into local components which commute with each other. We then construct an explicit equivalence with a block partitioned QCA.

### v2

@@ E.2 Covariant right restrictions
for all $g,k\in G$. The failure of the right restriction to be covariant is captured by local unitaries $\operatorname{\Psi}_{g}(k)$ which are uniquely defined up to phase by
@@
for all $g,k,l,m\in G$. They therefore define a twisted cohomology class $[\lambda]\in H^{2}(G,U(1)[G])$. (See Appendix G for the relevant definitions.)
The class $[\lambda]$ depends only on the symmetry $\alpha$, i.e. it does not depend on the choice of right restriction or the site $j$. So we obtain a well defined map
@@
In particular, the obstruction reduces to a homomorphism of monoids $\Lambda:(\operatorname{\mathsf{Sym}}_{G}/\sim)\rightarrow H^{2}(G,U(1)[G])$.
The proof can be found in Appendix I.
@@
    We will show in Appendix K that the obstruction $\Lambda(\alpha)$ is given by the inverse of the slant product of the anomaly of $\alpha$ (see Appendix G for definitions). In the context of SPTs the obstruction $\Lambda$ is therefore intimately related to the types and properties of the anyons supported by the gauged SPT [24], which are believed to be described by the Dijkgraaf-Witten TQFT [25] corresponding to the anomaly.
As the name suggests, if a symmetry $\alpha$ has vanishing obstruction to covariant right restrictions, then (after stacking with local degrees of freedom) $\alpha$ indeed admits covariant right restrictions:
@@
Let $\alpha$ be a symmetry of range $R$ such that $\Lambda(\alpha)=[1]$. For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by stacking with a local degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain $\widetilde{\mathcal{A}}$. Then the symmetry $\tilde{\alpha}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}$, where $\rho_{\operatorname{\mathrm{reg}}}$ is the left regular representation of $G$ on the local degree of freedom, admits a covariant right restriction at $j$ with defect size $5R$.
@@
$V_{g}^{*}\,\tilde{\alpha}^{(k)}\big{(}V_{\bar{k}gk}\big{)}$ $=\left(\sum_{l_{1}}\,\operatorname{\Psi}_{g}(l_{1})^{*}\otimes|l_{1}\rangle\langle l_{1}|\right)\times\left(\sum_{l_{2}}\,\tilde{\alpha}^{(k)}\big{(}\operatorname{\Psi}_{\bar{k}gk}(l_{2})\big{)}\otimes|kl_{2}\rangle\langle kl_{2}|\right)$
putting $l=l_{1}=kl_{2}$ and using Eq. (E.2) this becomes
@@
Now consider the right restriction $\tilde{\beta}_{\geq}$ of $\tilde{\alpha}$ with components $\tilde{\beta}_{\geq}^{(g)}=\mathrm{Ad}[V_{g}]\circ\tilde{\alpha}_{\geq}^{(g)}$. Then
@@
Suppose $\alpha$ has range $R$. We can stack the on-site algebra ${\mathcal{A}}_{j}$ by $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an extended spin chain $\widetilde{\mathcal{A}}$. By Lemma E.1 there is a right restriction $\tilde{\alpha}_{\geq}$ of $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ at $j$ with defect size $5R$ and such that $g\mapsto\tilde{\alpha}_{\geq}^{(g)}$ is a group homomorphism. We can therefore regard $\tilde{\alpha}_{\geq}$ as a symmetry.
Since $\tilde{\alpha}_{\geq}$ and $\alpha$ agree everywhere to the right of the site $j+5R$ it follows from local computability of the obstruction to covariant right restrictions (item 2 of Proposition E.2) that $\Lambda(\alpha)=\Lambda(\tilde{\alpha}_{\geq})$. But $\tilde{\alpha}_{\geq}$ agrees with $\mathrm{id}$ everywhere to the left of the site $j-5R$, so $\Lambda(\tilde{\alpha}_{\geq})=\Lambda(\mathrm{id})=[1]$ by local computability. We conclude that $\Lambda(\alpha)=[1]$. ∎
E.3 An invariant for symmetries that admit covariant right restrictions
Suppose $\alpha$ is a range $R$ symmetry such that $\Lambda(\alpha)=[1]$. Then, Lemma E.4 shows that, if we add a $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$-ancilla at any site $j$ and extend the symmetry to $\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}$, the new symmetry allows a covariant right restriction at $j$ with defect size $5R$.
For simplicity, let us assume that we have stacked with $\rho_{\operatorname{\mathrm{reg}}}$ everywhere so that $\alpha$ does allow covariant right restrictions of defect size $5R$ everywhere. Let $\alpha_{\geq}$ be such a covariant right restriction at some site $j\in{\mathbb{Z}}$, and let $\Phi(g,h)$ be fusion operators for this right restriction. Then we have
$\Phi(f,g)\Phi(fg,h)=\omega(f,g,h)\times\alpha_{\geq}^{(f)}\big{(}\Phi(g,h)\big{)}\Phi(f,gh)$
for a 3-cocycle $\omega$ with $[\omega]=\Omega(\alpha)$. Covariance of $\alpha_{\geq}$ applied to the defining property of the fusion operators
$\alpha_{\geq}^{(g)}\circ\alpha_{\geq}^{(h)}=\mathrm{Ad}[\Phi(g,h)]\circ\alpha_{\geq}^{(gh)}$ (E.3)
implies that there are phases $\mu_{g,h}(k)$ such that
$\alpha^{(k)}(\Phi(\bar{k}gk,\bar{k}hk))=\mu_{g,h}(k)\times\Phi(g,h).$ (E.4)
By straight computation, it follows that
$\alpha^{(kl)}\big{(}\Phi(\overline{kl}gkl,\overline{kl}hkl)\big{)}$ $=\mu_{g,h}(kl)\times\Phi(g,h)$
$=\alpha^{(k)}\big{(}\mu_{\bar{k}gk,\bar{k}hk}(l)\times\Phi(\bar{k}gk,\bar{k}hk)\big{)}=\mu_{g,h}(k)\mu_{\bar{k}gk,\bar{k}hk}(l)\times\Phi(g,h)$
hence
$\mu_{g,h}(k)\mu_{\bar{k}gk,\bar{k}hk}(l)=\mu_{g,h}(kl)$ (E.5)
for all $k,l,g,h\in G$. The phases $\mu_{g,h}(k)$ for $g,h,k\in G$ give rise to a map $k\to\mu(k)$ from $G$ to the $G$-module $U(1)[G^{2}]$ (We refer to Appendix G for the relevant definitions). Eq. (E.5) guarantees that such a map is a twisted 1-cocycle into $U(1)[G^{2}]$ (see (G.9)) and it has an associated class $[\mu]\in H^{1}(G,U(1)[G^{2}])$.
In general, such a class is not an invariant of the classification of locality preserving symmetries: under a different choice of covariant right-restriction $\beta^{(g)}_{\geq}=\mathrm{Ad}(U_{g})\circ\alpha_{\geq}^{(g)}$, the phases transform as
$\mu_{g,h}(k)\longrightarrow\frac{c_{g}(k)c_{h}(k)}{c_{gh}(k)}\mu_{g,h}(k)$ (E.6)
(see Appendix J for details), where the map $k\mapsto c(k)$ is a representative of a class $[c]\in H^{1}(G,U(1)[G])$. In Lemma E.9 we will prove that $[c]$ is not a topological obstruction, and can be made trivial by a local extension of $\alpha$. As a consequence, an invariant for symmetries that admit covariant right restrictions can be constructed by identifying classes in $H^{1}(G,U(1)[G^{2}])$ that differ by an element of the image of $H^{1}(G,U(1)[G])$ under a suitable homomorphism.
@@ E.3 An invariant for symmetries that admit covariant right restrictions
Let $\theta\in C^{1}(G,U(1)[G])$. Then $\iota(\theta)\in C^{1}(G,U(1)[G^{2}])$, pointwise defined by
$\iota(\theta)_{g,h}(k):=\frac{\theta_{g}(k)\theta_{h}(k)}{\theta_{gh}(k)}$
induces an injective group homomorphism $\iota:H^{1}(G,U(1)[G])\to H^{1}(G,U(1)[G^{2}])$. Hence the quotient
$\mathfrak{K}:=\frac{H^{1}(G,U(1)[G^{2}])}{\iota(H^{1}(G,U(1)[G]))}$ (E.7)
is a well-defined finite abelian group.
@@
Let $\theta\in C^{1}(G,U(1)[G])$ be a 1-cocycle from $G$ to $U(1)[G]$. That $\iota(\theta)$ is a 1-cocycle and that $\iota$ is a group homomorphism follow easily from the definitions. We prove injectivity, i.e., that $[\iota(\theta\delta\nu)]=[\iota(\theta)]\in H^{1}(G,U(1)[G^{2}])$ for a 1-coboundary $\delta\nu$ or, equivalently, that $[\iota(\delta\nu)]=[1]$, for $\nu\in U(1)[G]$. Firstly,
$(\delta\nu)_{g}(k)=\frac{k\cdot\nu_{g}}{\nu_{g}}=\frac{\nu_{\bar{k}gk}}{\nu_{g}},$
hence
$(\iota\circ\delta(\nu))_{g,h}(k)$ $=\frac{(\delta\nu)_{g}(k)(\delta\nu)_{h}(k)}{(\delta\nu)_{gh}(k)}=\frac{\nu_{\bar{k}gk}}{\nu_{g}}\frac{\nu_{\bar{k}hk}}{\nu_{h}}\frac{\nu_{gh}}{\nu_{\bar{k}ghk}}=\frac{(\iota(\nu))_{\bar{k}gk,\bar{k}hk}(k)}{(\iota(\nu))_{g,h}(k)}=(\delta\circ\iota(\nu))_{g,h}(k).$
We are then ready to define an invariant on the submonoid $\operatorname{\mathsf{Sym}}_{G}^{\Lambda=[1]}$ of $\operatorname{\mathsf{Sym}}_{G}$, which consists of all symmetries $\alpha$ for which $\Lambda(\alpha)=[1]$:
Proposition E.7.
The phases $\mu_{g,h}(k)$ satisfy the twisted 1-cocycle relation (E.5), hence they define a class $[\mu]\in\mathfrak{K}$, that depends only on the symmetry $\alpha$, i.e. it does not depend on the choice of right restriction or the site $j$. We obtain a well defined map
$\Upsilon:\operatorname{\mathsf{Sym}}_{G}^{\Lambda=[1]}$ $\longrightarrow\mathfrak{K}$
$\alpha$ $\longrightarrow\Upsilon(\alpha):=[\mu].$
Moreover, for symmetries $\alpha$ and $\beta$ of range $R$ this obstruction satisfies
  1. 1.
    If $\alpha$ is decoupled then $\Upsilon(\alpha)=[1]$, the identity element of $\mathfrak{K}$.
  2. 2.
    $\Upsilon$ is locally computable: If $\alpha$ and $\beta$ act on the same spin chain ${\mathcal{A}}$ and there is an interval $I$ of length $112R+$ 1 such that $\alpha|_{{\mathcal{A}}_{I}}=\beta|_{{\mathcal{A}}_{I}}$ then $\Upsilon(\alpha)=\Upsilon(\beta)$.
  3. 3.
    $\Upsilon$ is multiplicative under stacking: $\Upsilon(\alpha\otimes\beta)=\Upsilon(\alpha)\cdot\Upsilon(\beta)$.
  4. 4.
    $\Upsilon$ is constant on stable equivalence classes: $\alpha\sim\beta\,\implies\,\Upsilon(\alpha)=\Upsilon(\beta)$.
In particular, the map $\Upsilon$ reduces to a homomorphism of monoids $\Upsilon:{(\operatorname{\mathsf{Sym}}_{G}^{\Lambda=[1]}/\sim)}\to\mathfrak{K}.$
@@
See Appendix J. ∎
Lemma E.8.
If $\alpha$ has trivial anomaly, then also $\Lambda(\alpha)$ is trivial, so $\Upsilon(\alpha)$ is well defined. We have in this case $\Upsilon(\alpha)=[1]$.
Proof.
Same as the proof of Lemma E.5. ∎
E.4 Covariant right restrictions that are group homomorphisms
Lemma E.9.
Let $\alpha$ be a symmetry of range $R$ with trivial anomaly on a spin chain ${\mathcal{A}}$. For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by stacking with three local degrees of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})\otimes\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})\otimes\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain ${\mathcal{A}}^{\prime\prime\prime}$. There is an extension $\alpha^{\prime\prime\prime}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}^{\otimes 2}\otimes\rho_{\mathrm{ad}}$ of $\alpha$ to the enlarged spin chain and a covariant right restriction $\beta_{\geq}$ of $\alpha^{\prime\prime\prime}$ at $j$ with defect size $7R$ such that $g\mapsto\beta_{\geq}^{(g)}$ is a group homomorphism.
Proof.
Fix $j\in{\mathbb{Z}}$. We stack the on-site algebra ${\mathcal{A}}_{j}$ with a local degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain ${\mathcal{A}}^{\prime}$. Let $\alpha^{\prime}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}$ be the enlarged symmetry on ${\mathcal{A}}^{\prime}$. Then, Lemmas E.4 and E.5 imply that there is a covariant right restriction $\alpha^{\prime}_{\geq}$ of $\alpha^{\prime}$ at $j$ of defect size $5R$.
Since $\Lambda(\alpha^{\prime})=[1]$ we have a well defined invariant $\Upsilon(\alpha^{\prime})$, which is trivial by Lemma E.8. Recall that, by the defining Eq. (E.4),
$(\alpha^{\prime})^{(k)}\big{(}\Phi(\bar{k}gk,\bar{k}hk)\big{)}=\mu_{g,h}(k)\times\Phi(g,h),$
where $\Phi(g,h)$ are fusion operators for the covariant right restriction $\alpha^{\prime}_{\geq}$. Triviality of $\Upsilon(\alpha^{\prime})=[\mu]$ implies
$\mu_{g,h}(k)=\frac{\nu_{\bar{k}gk,\bar{k}hk}}{\nu_{g,h}}\frac{c_{g}(k)c_{h}(k)}{c_{gh}(k)}$
for $U(1)$ phases $\nu_{a,b}$ and a representative $c$ of a class $[c]\in H^{1}(G,U(1)[G])$. Without loss, we choose new fusion operators $\nu_{g,h}\Phi(g,h)$, and denote them again by $\Phi(g,h)$.
Further stacking by another local degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$ at site $j$, we obtain an enlarged spin chain ${\mathcal{A}}^{\prime\prime}$. We extend the symmetry to $\alpha^{\prime\prime}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}^{\otimes 2}$ and take a right restriction $\alpha^{\prime\prime}_{\geq}$ with components $(\alpha_{\geq}^{\prime\prime})^{(g)}=(\alpha^{\prime}_{\geq})^{(g)}\otimes\mathrm{Ad}[U(g)]$ with
$U(g)=\sum_{l\in G}c_{g}(l)|l\rangle\langle l|.$
By the twisted 1-cocycle relation (G.4) satisfied by $c$, one easily checks that $\rho_{\operatorname{\mathrm{reg}}}^{(k)}(U(\bar{k}gk))=\overline{c_{g}(k)}U(g)$. It follows that $\alpha^{\prime\prime}_{\geq}$ is a covariant right restriction of $\alpha^{\prime\prime}$, with fusion operators
$\Phi^{\prime\prime}(g,h)=\Phi(g,h)\otimes U(g)U(h)U(gh)^{*},$
from which we compute
$(\alpha^{\prime\prime})^{(k)}\big{(}\Phi^{\prime\prime}(\bar{k}gk,\bar{k}hk)\big{)}=\frac{c_{gh}(k)}{c_{g}(k)c_{h}(k)}\mu_{g,h}(k)\times\Phi^{\prime\prime}(g,h)=\Phi^{\prime\prime}(g,h).$
We now add another local degree of freedom $\operatorname{\mathrm{End}}{{\mathbb{C}}^{\left|G\right|}}$ at site $j$ and, as in the proof of Lemma E.1, we define unitaries $V(g)=\sum_{l}\operatorname{\Phi}^{\prime\prime}(g,l)\otimes|l\rangle\langle gl|$ supported on $[j-5R,j+6R]$ so that $\beta_{\geq}^{(g)}=\mathrm{Ad}[V(g)^{*}]\circ(\alpha^{\prime\prime}_{\geq}\otimes\mathrm{id})$ defines a right restriction of $\alpha^{\prime\prime\prime}=\alpha^{\prime\prime}\otimes\rho_{\mathrm{ad}}$, where $g\mapsto\beta_{\geq}^{(g)}$ is a group homomorphism. Here $\rho_{\mathrm{ad}}^{(k)}=\mathrm{Ad}[J^{(k)}]$ with $J^{(k)}|h\rangle=|kh\bar{k}\rangle$.
Moreover,
$(\alpha^{\prime\prime\prime})^{(k)}\big{(}V(\bar{k}gk)\big{)}=\sum_{l\in G}\operatorname{\Phi}^{\prime\prime}(g,l)\otimes|l\rangle\langle gl|=V(g).$
This implies that $\beta_{\geq}^{(g)}=\mathrm{Ad}[V(g)^{*}]\circ(\alpha^{\prime\prime}_{\geq}\otimes\mathrm{id})$ is a covariant right restriction of $\alpha^{\prime\prime\prime}$ at $j$ with defect size $7R$ which is itself a group morphism. ∎
Appendix F Proof of the main Theorem B.1
In this section we prove injectivity of the map $\Omega:\operatorname{\mathsf{Sym}}_{G}\rightarrow H^{3}(G,U(1))$ (Corollary F.3), and prove the main classification Theorem B.1.
F.1 Injectivity of $\Omega$
Firstly, we prove that a symmetry with trivial anomaly is stably equivalent to a symmetry that can be written as a product of mutually commuting local representations:
Lemma F.1.
Let $\alpha$ be a symmetry on a spin chain ${\mathcal{A}}$, with trivial anomaly. Then $\alpha\sim\alpha^{\prime\prime}$, where $\alpha^{\prime\prime}$ is a symmetry that can be written as a formal product
$(\alpha^{\prime\prime})^{(g)}=\prod_{j\in{\mathbb{Z}}}(\alpha^{\prime\prime})_{j}^{(g)},\qquad(\alpha^{\prime\prime})^{(g)}_{j}=\mathrm{Ad}(\tilde{U}_{j}(g)),$ (F.1)
where
  1. 1.
    $g\mapsto(\alpha_{j}^{\prime\prime})^{(g)}$ is a group homomorphism,
  2. 2.
    $(\alpha^{\prime\prime})^{(g)}_{j}$ is supported on $[jL-7R,(j+1)L+8R]$ for all $g\in G$,
  3. 3.
    $[\tilde{U}_{i}(g),\tilde{U}_{j}(h)]=0$, whenever $i\neq j$, for all $g,h\in G$.
Proof.